Graph the system of equations on graph paper.
{8x+6y=482x−3y=−6
Which statements are true about the solution to the system of equations?
Select each correct answer.


The ordered pair that is the solution to the system lies in Quadrant I .
The ordered pair that is the solution to the system lies in Quadrant II .
The x-coordinate of the solution is −3 .
The y-coordinate of the solution is 3.
The x-coordinate of the solution is 3.
The y-coordinate of the solution is 4.

Answer by Theo(13342)

your equations are:

8x + 6y = 48
2x - 3y = -6

multiply both sides of the second eqution by 4 and leave the first equation as is to get:

8x + 6y = 48
8x - 12y = -24

subtract the second equation from the first to get:

18y = 72

solve for y to get y = 4

go back to the first original equation and solve for x to get x = 3.

your solution is x = 3 and y = 4

answers to your questions are:

The ordered pair that is the solution to the system lies in Quadrant I . TRUE
The ordered pair that is the solution to the system lies in Quadrant II . FALSE
The x-coordinate of the solution is −3 . FALSE
The y-coordinate of the solution is 3. FALSE
The x-coordinate of the solution is 3. TRUE
The y-coordinate of the solution is 4. TRUE

the graph of the these equations and their intersection is shown below.

$$$